OpenStudy (hannaha):

Medal and fan!! Evaluate for a = 2, b = 3, and c = 4. 3a2 + 4b2

OpenStudy (anonymous):

3(2)(2) + 4(3)(2) = 12+24 =36

OpenStudy (hannaha):

Tried that, it's wrong :/

OpenStudy (hannaha):

It looks like the answer is 48...but I have no idea how to get there

OpenStudy (deadpoolroxz):

whats dah question

OpenStudy (hannaha):

Evaluate for a = 2, b = 3, and c = 4. 3a^2 + 4b^2= ?

OpenStudy (deadpoolroxz):

btw ever thot of using a calculator...

OpenStudy (hannaha):

Oh believe me I have...

OpenStudy (deadpoolroxz):

(3x(2^2))+(4x(3^2))

OpenStudy (deadpoolroxz):

pemdas

OpenStudy (deadpoolroxz):

sorry i dont have me calc on me right now hope it some wat helps

OpenStudy (hannaha):

AHHHHHHHHH finally!!! THANK YOU

OpenStudy (deadpoolroxz):

okie medal me ill medal u

OpenStudy (anonymous):

12 + 36 = 48

OpenStudy (deadpoolroxz):

2 l8te bruh

OpenStudy (hannaha):

lol

OpenStudy (hannaha):

thanks though @robtobey

OpenStudy (deadpoolroxz):

hannah wana chat

OpenStudy (hannaha):

sure, i'll be doing homework too tho, or trying to :D

OpenStudy (deadpoolroxz):

go 2 mathmatics @ the bottom